OpenStudy (anonymous):

factor : 2x^2+5x-7 and 2x^2+15x+28

OpenStudy (anonymous):

2x^2+5x-7 (2x+7)(x-1)

OpenStudy (anonymous):

2x^2+15x+28 (2x+7)(x+4)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[2x ^{2} + 5x - 7\]\[x ^{2} + 5x - 14\]\[(x - 7) (x + 2)\]\[x _{1} = \frac{ 7 }{ 2 } = 3\frac{ 1 }{ 2 }\]\[x = -\frac{ 2 }{ 2 } = 1\]

OpenStudy (anonymous):

\[2x ^{2} + 15x + 28\]\[x ^{2} + 15x + 56\]\[(x + 8) (x + 7)\]\[x _{1} = - \frac{ 8 }{ 2 } = -4\]\[x = - \frac{ 7 }{ 2 } = -3 \frac{ 1 }{ 2 }\]